OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = -1/5 and point (5, -2). A. y + 2 = -(1/5)(x - 5) B. y + 5 = -(1/5)(x - 2) C. y – 5 = (1/5)(x + 2) D. y – 5 = (1/5)(x - 2)

jimthompson5910 (jim_thompson5910):

Similar Problem: Write an equation in point slope form given: slope, m, = -1/3 and point (7, -3). y - y1 = m(x - x1) y - y1 = (-1/3)(x - x1)... plug in the slope m = -1/3 y - (-3) = (-1/3)(x - 7)... plug in the given point (x1,y1) = (7,-3) y + 3 = (-1/3)(x - 7)